Skip to main content

AAVSOnet status, 2018-07-04

1 post / 0 new
HQA's picture
AAVSOnet status, 2018-07-04

Here is the current system status.

BSM_Berry  operational

BSM_Hamren  camera problem, with vertical streaking.  There is a good chance that this is a power supply problem, so we are sending a replacement power supply to Hawaii.  If this is successful, BSM_Hamren should be back into operation next week.

BSM_NH  operational.  We should begin tests with the CMOS camera in mid-July.

BSM_NM  into summer monsoon.  Little or no data expected before mid-September.

BSM_S  operational, now that Peter solved the focuser problem by replacing a cable.  However, the system is off-line for a while as Peter is on holiday.

OC61  off-line until at least mid-July.  The dome motion motor failed, and the replacement motor is not a perfect fit.  Nigel needs to do some machining, and he is on travel.

SRO  operational, but the summer monsoon is starting.  Little or no data expected before mid-September.

TMO61  various issues that are not expected to be resolved until mid-July.

AAVSO 49 Bay State Rd. Cambridge, MA 02138 617-354-0484